The spot rate between Canada and the U.S. is Can$1.2398/$, while the one-year forward rate is Can$1.2397/$. The risk-free rate in Canada is 4.31 percent and risk-free rate in the United States is 2.60 percent. How much in profit can you earn on $6,500 utilizing covered interest arbitrage?
Answer #1

Step 1 Borrow 6500$ and convert into Can$ at Spot Rate

6500 borrowed at interest rate of 2.60% and converted to Can$ i.e. 6500 * 1.2398 = Can $ 8058.7

Step 2 invest Can$ 8058.7 in Canada @ 4.31% and receive interest

8058.7 + 8058.7*4.31% = Can$8406.03

Step 3 convert Back to $ at Forward Rate and pay the interest

Can$8406.03 / 1.2397 - (6500 * 2.60%) = $ 6611.69

Arbitrage profit will be = 6611.69 - 6500 = $ 111.70